# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ Abstract
Graph generation is fundamental in diverse scientific applications, due to its ability to reveal the underlying distribution of complex data, and eventually generate new, realistic data points. Despite the success of diffusion models in this domain, those face limitations in sampling efficiency and flexibility, stemming from the tight coupling between the training and sampling stages. To address this, we propose DeFoG, a novel framework using discrete flow matching for graph generation. DeFoG employs a flow-based approach that features an efficient linear interpolation noising process and a flexible denoising process based on a continuous-time Markov chain formulation. We leverage an expressive graph transformer and ensure desirable node permutation properties to respect graph symmetry. Crucially, our framework enables a disentangled design of the training and sampling stages, enabling more effective and efficient optimization of model performance. We navigate this design space by introducing several algorithmic improvements that boost the model performance, consistently surpassing existing diffusion models. We also theoretically demonstrate that, for general discrete data, discrete flow models can faithfully replicate the ground truth distribution - a result that naturally extends to graph data and reinforces DeFoG’s foundations. Extensive experiments show that DeFoG achieves state-of-the-art results on synthetic and molecular datasets, improving both training and sampling efficiency over diffusion models, and excels in conditional generation on a digital pathology dataset.
@@ 1 Introduction
Graph generation has become a fundamental task across diverse fields, from molecular chemistry to social network analysis, due to graphs’ capacity to represent complex relationships and generate novel, realistic structured data. Diffusion-based graph generative models (Niu et al., 2020; Jo et al., 2022), particularly those tailored for discrete data to capture the inherent discreteness of graphs (Vignac et al., 2022; Siraudin et al., 2024), have emerged as a compelling approach, demonstrating state-of-the-art performance in applications such as molecular generation (Irwin et al., 2024), reaction pathway design (Igashov et al., 2024), neural architecture search (Asthana et al., 2024), and combinatorial optimization (Sun & Yang, 2024). However, the sampling processes of these methods remain highly constrained due to a strong interdependence with training: design choices made during training mostly determine the options available during sampling. Therefore, optimizing parameters such as noise schedules requires re-training, which incurs significant computational costs.
Recently, Campbell et al. (2024); Gat et al. (2024) introduced a discrete flow matching (DFM) framework, which maps noisy to clean discrete data with iterations similar to diffusion models. However, DFM has demonstrated superior performance compared to discrete diffusion models by leveraging a straightforward and efficient noising process, and a flexible formulation based on continuous-time Markov chain (CTMC) for denoising. Specifically, its noising process involves a linear interpolation in the probability space between noisy and clean data, displaying well-established advantages for continuous state spaces in terms of performance (Esser et al., 2024; Lipman et al., 2023) and theoretical properties (Liu et al., 2023). Most notably, the denoising process in DFM uses an adaptable sampling step scheme (both in size and number) and CTMC rate matrices that are independent of the training setup. Such a decoupling allows for the design of sampling algorithms that are disentangled from training, enabling efficient performance optimization at the sampling stage without extensive retraining. However, despite its promising results, the applicability of DFM to graph generation remains unclear, and the proper exploitation of its additional training-sampling flexibility is still an open question in graph settings.
In this work, we propose DeFoG, a novel and versatile DFM framework for graph generation that operates under theoretical guarantees, featuring a linear interpolation noising process and CTMC-based denoising process inherited from the DFM model. Our contribution begins by proposing the first discrete flow matching model tailored to graphs, ensuring node permutation equivariance to respect graph symmetry and addressing the model expressivity limitations inherent to this data modality (Morris et al., 2018). Additionally, to “defog” the design space of this decoupled training-sampling framework, we extend and propose novel training and sampling algorithms, including alternative initial distributions, modifications of the CTMC rate matrices that govern the denoising trajectory, and time-adaptive strategies designed to align with dataset-specific characteristics. We explicitly demonstrate that these algorithmic improvements, which can be seamlessly integrated, accelerate convergence and enhance generation performance by a large margin. To assert DeFoG’s strong theoretical foundation, we extend existing guarantees of DFM models for general discrete data. First, we reinforce the pertinence of the loss function by explicitly relating it to the model estimation error. Then, we demonstrate that the distribution generated by the denoising process in DFM models remains faithful to the ground truth distribution. Both theorems extend naturally to graph data, providing strong support for the design choices of our new graph generative framework.
Our experiments show that DeFoG achieves state-of-the-art performance on both synthetic graph datasets with diverse topologies and complex molecular datasets. Notably, while the vanilla DeFoG framework performs on par with diffusion models, our tailored training-sampling optimization pipeline permits DeFoG to consistently outperform these models. Furthermore, DeFoG achieves performance comparable to certain diffusion models with only 5 to 10% of the sampling steps, highlighting significant efficiency gains. We also underscore DeFoG’s versatility through a conditional generation task in digital pathology, where it notably outperforms existing unconstrained models. In a nutshell, DeFoG opens the door to more flexible graph generative models that can efficiently handle a wide range of tasks in diverse domains with improved performance and reduced computational costs.

### v3

@@ Abstract
Graph generative models are essential across diverse scientific domains by capturing complex distributions over relational data. Among them, graph diffusion models achieve superior performance but face inefficient sampling and limited flexibility due to the tight coupling between training and sampling stages. We introduce DeFoG, a novel graph generative framework that disentangles sampling from training, enabling a broader design space for more effective and efficient model optimization. DeFoG employs a discrete flow-matching formulation that respects the inherent symmetries of graphs. We theoretically ground this disentangled formulation by explicitly relating the training loss to the sampling algorithm and showing that DeFoG faithfully replicates the ground truth graph distribution. Building on these foundations, we thoroughly investigate DeFoG’s design space and propose novel sampling methods that significantly enhance performance and reduce the required number of refinement steps. Extensive experiments demonstrate state-of-the-art performance across synthetic, molecular, and digital pathology datasets, covering both unconditional and conditional generation settings. It also outperforms most diffusion-based models with just 5–10% of their sampling steps.
Graph Generative Models, Discrete Flow Matching
[…]
@@ 1 Introduction
Graph generation has become a fundamental task across diverse fields, from molecular chemistry to social network analysis, due to graphs’ capacity to represent complex relationships and generate realistic structured data. Diffusion-based graph generative models (Niu et al., 2020; Jo et al., 2022), particularly those tailored for discrete data (Vignac et al., 2022), have emerged as compelling approaches, demonstrating pioneering performance in applications such as molecular generation (Irwin et al., 2024), reaction pathway design (Igashov et al., 2024), neural architecture search (Asthana et al., 2024), and combinatorial optimization (Sun & Yang, 2024). Recently, continuous-time discrete diffusion frameworks have further advanced the domain of discrete graph diffusion (Xu et al., 2024; Siraudin et al., 2024). These frameworks leverage the robustness and flexibility of continuous-time modeling, while preserving the natural alignment with the discrete structure of graphs.
Despite their state-of-the-art performance, the training and sampling stages of diffusion-based models is tightly entangled, restricting sampling options to training-phase choices. Thus, optimizing components such as noise schedules or rate matrices requires re-training for each configuration, resulting in prohibitive computational costs. Consequently, these models often adopt a single configuration across graph datasets. This one-size-fits-all approach fails to accommodate the diverse structural characteristics of different datasets, leaving room for further improvement.
(a) Motivation of DeFoG.
(b) Overview of DeFoG.
Figure 2: (a) DeFoG enhances graph generation by introducing training-sampling decoupling, an orthogonal improvement within graph iterative refinement models, while preserving the sampling flexibility and inherent discreteness exploitation of prior SOTA models. (b) One node, $x^{(n)}$, is selected to illustrate both noising and denoising processes. For noising, DeFoG follows a straight path from the one-hot encoding $p_{1}$ of the clean node to the initial distribution $p_{0}$. For denoising, a network parameterized by $\theta$ predicts the marginal distributions of the clean graph, there the node’s distribution ${\bm{p}^{\theta,(n)}_{1|t}(\cdot|G_{t})}$ is used to compute its rate matrix $\bm{R}_{t}^{\theta,(n)}$ and, subsequently, its probability at the next time point $t+\Delta t$.
In this work, we present DeFoG, a novel graph generative framework that disentangles the training and sampling stages (Figure 2(a)), addressing the inefficiencies in graph diffusion models and achieving state-of-the-art (SOTA) performance. DeFoG leverages a discrete flow matching (DFM) inspired formulation (Campbell et al., 2024) that we tailor to graph settings. It features a linear interpolation noising process and a continuous-time Markov chain (CTMC)-based denoising process, while ensuring node permutation equivariance and addressing the model expressivity limitations inherent to this data modality (Morris et al., 2019). We demonstrate that training-sampling decoupling not only enhances flexibility but is also provably sound. By theoretically establishing that training loss optimization leads to improved sampling dynamics, DeFoG enables faithful replication of the ground truth graph distribution. To navigate the expanded design space enabled by such disentanglement, we take a critical step by “defogging” this space. Specifically, we explore and propose various sampling methods, including time-adaptive methods and modifications to CTMC rate matrices, to better govern denoising trajectories and align with the unique characteristics of graph datasets.
Our experiments show that DeFoG achieves SOTA performance across diverse datasets, with near-saturated validity of 99.5%, 96.5%, and 90% on planar, tree, and stochastic block model (SBM) datasets, respectively. On complex molecular data, it achieves 92.8% validity on MOSES (Polykovskiy et al., 2020), surpassing the previous SOTA of 90.5%. Moreover, DeFoG achieves 95.0% and 86.5% validity on planar and SBM datasets, respectively, with only 5–10% of the sampling steps used by diffusion models. This performance surpasses all but one diffusion model on the planar dataset and ranks best on SBM, highlighting substantial efficiency gains. To further highlight the versatility of DeFoG, we also test it in conditional generation tasks for digital pathology, where it largely outperforms existing unconstrained models. Ablation studies further confirm the need of each proposed sampling method and highlight the importance of dataset-specific sampling procedures to effectively address diverse data characteristics.
Our main contributions are as follows:
  • We introduce DeFoG, a novel graph generative model that effectively exploits the training-sampling decoupling enabled by its flow-based formulation, significantly enhancing sampling flexibility and efficiency.
  • We provide a theoretical foundation for both our training and sampling algorithms, validating the soundness of the disentanglement framework;
  • We comprehensively explore DeFoG’s design space with novel training and sampling approaches, highlight critical configurations for graph data and attain state-of-the-art performance across diverse datasets.
Overall, DeFoG enables more effective graph generation with reduced computational costs under theoretical guarantees, paving the way for broader adoption of graph generative models in real-world applications.
2 Background
In generative modeling, the primary goal is to generate new data samples from the underlying distribution that produced the original data, $\bm{p}_{\mathrm{data}}$. An effective approach is to learn a mapping between a simpler distribution $\bm{p}_{\epsilon}$ that can be easily sampled, and $\bm{p}_{\mathrm{data}}$.
Iterative refinement models achieve this mapping through a stochastic process over the time interval $t\in[0,1]$ for variables in discrete state spaces. For the sake of simplicity, we describe here an univariate formulation. At any time $t$, we consider a discrete variable with $Z$ possible values, denoted by $z_{t}\in\mathcal{Z}=\{1,\ldots,Z\}$. The marginal distribution of $z_{t}$ is represented by the vector $\bm{p}_{t}\in\Delta^{Z}$, with $\Delta^{Z}=\left\{\mathbf{u}\in\mathbb{R}^{Z}\mid\sum_{i=1}^{Z}u_{i}=1,\;u_{i}\geq 0,\,\forall i\right\}$. The initial distribution is set to a predefined noise distribution, $\bm{p}_{0}=\bm{p}_{\epsilon}$, while $\bm{p}_{1}=\bm{p}_{\text{data}}$ represents the target data distribution. We refer to the mapping $t:1\to 0$ as the noising process and $t:0\to 1$ as the denoising process.
DFM (Campbell et al., 2024) builds upon a streamlined noising process. In particular, the noising trajectory $p_{t|1}(z_{t}|z_{1})\in[0,1]$ is defined through a simple linear interpolation starting from a chosen datapoint $z_{1}$:
$p_{t|1}(z_{t}|z_{1})=t\,\delta(z_{t},z_{1})+(1-t)\,p_{0}(z_{t}),$ (1)
where $\delta(z_{t},z_{1})$ is the Kronecker delta (1 when $z_{t}=z_{1}$). A usual choice for the initial distribution is the uniform distribution over the state space, $\bm{p}_{0}=[1/Z,\ldots,1/Z]$.
In the denoising stage, DFM leverages a CTMC formulation. In general, a CTMC is characterized by an initial distribution, $\bm{p}_{0}$, and a rate matrix, $\bm{R}_{t}\in\mathbb{R}^{Z\times Z}$ that governs its evolution across time $t\in[0,1]$. Specifically, the rate matrix defines the instantaneous transition rates between states, such that:
$p_{t+\mathrm{d}t|t}(z_{t+\mathrm{d}t}|z_{t})=\delta(z_{t},z_{t+\mathrm{d}t})+R_{t}(z_{t},z_{t+\mathrm{d}t})\textrm{d}t,$ (2)
where $R_{t}(z_{t},z_{t+\mathrm{d}t})$ denotes an entry in the rate matrix. Intuitively, $R_{t}(z_{t},z_{t+\mathrm{d}t})\mathrm{d}t$ yields the probability that a transition from state $z_{t}$ to state $z_{t+\mathrm{d}t}$ will occur in the next infinitesimal time step $\mathrm{d}t$. By definition, we have $R_{t}(z_{t},z_{t+\mathrm{d}t})\geq 0$ for $z_{t}\neq z_{t+\mathrm{d}t}$. Consequently, we further have $R_{t}(z_{t},z_{t})=-\sum_{z_{t+\mathrm{d}t}\neq z_{t}}R_{t}(z_{t},z_{t+\mathrm{d}t})$ to ensure normalization $\sum_{z_{t+\mathrm{d}t}}p_{t+\mathrm{d}t|t}(z_{t+\mathrm{d}t}|z_{t})=1$. Under this definition, the marginal distribution and the rate matrix of a CTCM are related by a conservation law, the Kolmogorov equation, given by $\partial_{t}\bm{p}_{t}=\bm{R}^{\top}_{t}\bm{p}_{t}$. If expanded, this expression unveils the time derivative of the marginal distribution as the net balance between the inflow and outflow of probability mass at that state.
Similarly to the noising process of Eq. 1, the denoising is also performed under conditioning on $z_{1}$. Specifically, we consider a $z_{1}$-conditional rate matrix, $\bm{R}_{t}(\cdot\,,\cdot|z_{1})\in\mathbb{R}^{Z\times Z}$, that will govern the denoising in DFM. Under mild assumptions, Campbell et al. (2024) present a closed-form for a valid conditional rate matrix, i.e., a matrix that verifies the corresponding Kolmogorov equation, for $z_{t}\neq z_{t+\mathrm{d}t}$, defined as:
$R^{*}_{t}{(z_{t},z_{t+\mathrm{d}t}|z_{1})}=\frac{\operatorname{ReLU}[\partial_{t}p_{t|1}(z_{t+\text{d}t}|z_{1})-\partial_{t}p_{t|1}(z_{t}|z_{1})]}{Z^{>0}_{t}\;p_{t|1}(z_{t}|z_{1})}$ (3)
and $Z^{>0}_{t}=|\{z_{t}:p_{t|1}(z_{t}|z_{1})>0\}|$. Again, normalization is performed for the case $z_{t+\mathrm{d}t}=z_{t}$. Intuitively, $R^{*}_{t}$ applies a positive rate to states needing more mass than the current state $z_{t}$ (details in Sec. B.5). Finally, it can be shown that $R_{t}(z_{t},z_{t+\mathrm{d}t})=\mathbb{E}_{p_{1|t}(z_{1}|z_{t})}[R_{t}(z_{t+\mathrm{d}t},z_{t}|z_{1})]$, which is employed in Eq. 2 for denoising.
While the DFM paradigm enables training-sampling disentanglement, it lacks a complete formulation and empirical validation on graph data. Moreover, how to effectively leverage this disentanglement to enhance sampling performance remains underexplored, particularly for graph-specific tasks. We introduce DeFoG to address these gaps.
3 DeFoG Framework
In this section, we present DeFoG (Discrete Flow Matching on Graphs), a novel iterative refinement framework for graph generation that leverages the decoupling of training and sampling stages. We begin by describing its noising and denoising processes, highlighting how they enable this disentanglement, as illustrated in Figure 2(b). We theoretically demonstrate that this flexible framework is also robust by proving that optimizing the training loss improves sampling dynamics, ensuring that DeFoG can faithfully replicate graph distributions. Then, we discuss the expanded design space enabled by DeFoG’s disentanglement, which drives key improvements in graph generation performance. Finally, we establish the node permutation equivariance/invariance guarantees of DeFoG.
3.1 Learning Discrete Flows over Graphs
We instantiate undirected graphs with $N$ nodes as $G=(x^{1:n:N},e^{1:i<j:N})$, where $x^{1:n:N}=(x^{(n)})_{1\leq n\leq N}$ and $e^{1:i<j:N}=(e^{(ij)})_{1\leq i<j\leq N}$ denote the node and edge sets, respectively, with $x^{(n)}\in\mathcal{X}=\{1,\dots,X\}$ and $e^{(ij)}\in\mathcal{E}=\{1,\dots,E\}$. Following standard practice in the field (Vignac et al., 2022; Xu et al., 2024; Siraudin et al., 2024), we consider an edge between every pair of nodes, where one of the edge categories explicitly represents the absence of an edge (i.e., a “non-existing” edge).)
Noising
We now define the noising process of DeFoG. According to Eq. 1, with shared initial distributions across nodes and edges, denoted as $p^{\mathcal{X}}_{0}$ and $p^{\mathcal{E}}_{0}$, respectively, we formulate the noising trajectory by independently adding noise to each node and each edge:
$p_{t|1}(G_{t}|G_{1})=\prod_{n}p_{t|1}\left(x^{(n)}_{t}|x^{(n)}_{1}\right)\prod_{i<j}p_{t|1}\left(e^{(ij)}_{t}|e^{(ij)}_{1}\right).$
Different $p^{\mathcal{X}}_{0}$ and $p^{\mathcal{E}}_{0}$ are further discussed in Sec. C.1.
Sampling
As formulated in Sec. 2, the denoising process requires simulating a CTMC, driven by its rate matrix $\bm{R}_{t}$. We start by sampling a purely noisy graph $G_{0}$ from the predefined initial distribution $p_{0}(G_{0})=\prod_{n}p^{\mathcal{X}}_{0}(x_{0}^{(n)})\prod_{i<j}p^{\mathcal{E}}_{0}(e^{ij}_{0})$. Then, we progress in the denoising process by employing independent Euler steps for each node and edge, with a finite time step $\Delta t$, i.e., we iteratively sample progressively denoised graphs from $\tilde{p}_{t+\Delta t|t}(G_{t+\Delta t}|G_{t})$, given by:
$\prod_{n}\tilde{p}^{(n)}_{t+\Delta t|t}(x^{(n)}_{t+\Delta t}|G_{t})\prod_{i<j}\tilde{p}^{(ij)}_{t+\Delta t|t}(e^{(ij)}_{t+\Delta t}|G_{t}).$ (4)
Each term $\tilde{p}^{(n)}_{t+\Delta t|t}(x^{(n)}_{t+\Delta t}|G_{t})$ corresponds to the Euler step given in Eq. 2, where the transition dynamics are governed by the rate matrix computed as:
[…]
and similarly for $\tilde{p}^{(ij)}_{t+\Delta t|t}(e^{(ij)}_{t+\Delta t}|G_{t})$.
Algorithm 1 DeFoG Training
[…]
Algorithm 2 DeFoG Sampling
[…]
Training
The rate matrix used in the denoising steps above requires the knowledge of the marginal distributions $\bm{p}^{(n)}_{{1|t}}(\cdot|G_{t})\in\Delta^{X}$ and $\bm{p}^{(ij)}_{{1|t}}(\cdot|G_{t})\in\Delta^{E}$ for all nodes and all edges, respectively. Both are gathered in $\bm{p}_{{1|t}}(\cdot|G_{t})=\left(\bm{p}^{1:n:N}_{{1|t}}(\cdot|G_{t}),\,\bm{p}^{1:i<j:N}_{{1|t}}(\cdot|G_{t})\right)$. Each of these components consists of the clear marginal distribution prediction given a noisy graph $G_{t}$. However, the computation of these terms is generally intractable. Instead, we train a neural network, parameterized by $\theta$, to approximate them, $\bm{p}^{\theta}_{{1|t}}(\cdot|G_{t})$. To cover different noise levels, we sample $t\sim\mathcal{T}$, where $\mathcal{T}$ is an arbitrary distribution over $[0,1]$. In DFM, $\mathcal{T}$ is typically set to the uniform distribution, though alternative choices can enhance performance, as later explored in Sec. 3.2. The training loss is naturally formulated as:
$\mathcal{L}_{\text{DeFoG}}=\mathbb{E}_{t\sim\mathcal{T},p_{1}(G_{1}),p_{t|1}(G_{t}|G_{1})}\operatorname{CE}_{\lambda}(G_{1},\bm{p}^{\theta}_{1|t}(\cdot|G_{t})),$
where $\operatorname{CE}_{\lambda}(G_{1},\bm{p}^{\theta}_{1|t}(\cdot|G_{t}))$ is defined as:
[…]
Here, $\lambda\in\mathbb{R}^{+}$ is introduced to weight nodes and edge differently to more flexibly capture varying topologies.
Decoupled Training and Sampling
DeFoG exhibits a clear disentanglement of training and sampling. The training phase focuses on predicting the marginal probabilities of the clean graph $\bm{p}^{\theta}_{1|t}(\cdot|G_{t})$, while sampling relies on the rate matrix formulation. Importantly, the training process is agnostic to the choice of the $z_{1}$-conditional rate matrix. Thus, different $z_{1}$-conditional rate matrices can be employed at sampling time, such as $\bm{R}^{\star}(\cdot,\cdot|z_{1})$ in Eq. 3. This decoupling of sampling from training provides additional flexibility in DeFoG’s design, which can be leveraged to further enhance performance. Notably, we further demonstrate that, upon this decoupling, optimizing DeFoG’s training loss improves its sampling dynamics, ensuring the soundness of our framework.
Corollary 1 (Bounded estimation error of rate matrix for graphs).
Given $t\in[0,1]$ and graphs $G_{t},G_{t+\mathrm{d}t}$, and $G_{1}\sim p_{1}(G_{1})$, there exist constants $\bar{C}_{0},\bar{C}_{1},\bar{C}_{3}>0$ such that the rate matrix estimation error can be upper bounded by:
$|R_{t}(G_{t},G_{t+\mathrm{d}t})-R^{\theta}_{t}(G_{t},G_{t+\mathrm{d}t})|^{2}\leq\bar{C}_{0}+$
$+\bar{C}_{1}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{n}-\log p^{\theta,{(n)}}_{1|t}(x^{(n)}_{1}|G_{t})\right]$
$+\bar{C}_{2}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{i<j}-\log p^{\theta,{(ij)}}_{1|t}(e^{(ij)}_{1}|G_{t})\right].$
By taking the expectation over $t\sim\mathcal{T}$ and summing over $G_{t}$, minimizing the derived upper bound of rate matrix estimation error in Corollary 1 with respect to $\theta$ corresponds directly to minimizing the loss function of DeFoG with $\lambda=1$. Therefore, we guarantee that our training loss minimization is aligned with accurate rate matrix estimation.
Upon this result, we are now in conditions of justifying DeFoG’s approximated denoising algorithm.
Corollary 2 (Bounded deviation of the generated graph distribution).
Let $p_{1}$ be the marginal distribution at $t=1$ of a groundtruth CTMC, $\{G_{t}\}_{0\leq t\leq 1}$, and $\tilde{p}_{1}$ be the marginal distribution at $t=1$ of its independent-dimensional Euler sampling approximation, with a maximum step size $\Delta t$. Then, under 6, the following total variation bound holds:
$\|p(G_{1})$ $-p_{\text{data}}\|_{\text{TV}}\leq\bar{U}\left(XN+E\frac{N(N-1)}{2}\right)\;$
$+\bar{B}\left(XN+E\frac{N(N-1)}{2}\right)^{2}{\Delta t}+\;O(\Delta t),$
where $\bar{U}$ and $\bar{B}$ are constant upper bounds for the bound from Corollary 1 and for the denoising process relative to its noising counterpart, respectively, for any $t\in[0,1]$.
The first term of the bound captures the estimation error introduced by using the neural network approximation $R^{\theta}_{t}(G_{t},G_{t+\mathrm{d}t})$. From Corollary 1, this term is bounded. The remaining terms arise from the discretization of the CTMC and can be controlled by reducing the step size $\Delta t$. Consequently, the deviation introduced by this approximation can be made arbitrarily small, ensuring that the generated distribution remains close to the groundtruth and validating our graph sampling scheme.
The resulting training and sampling processes are detailed in Algs. 1 and 2. The proofs of Corollaries 1 and 2 are provided in Sec. D.1.
3.2 Design Space of DeFoG
As described in the previous section, DeFoG benefits from greater flexibility due to its training-sampling decoupling. For example, it allows the number and size of sampling steps to be adjusted dynamically, unlike the fixed steps in discrete-time diffusion (Vignac et al., 2022), and enables adjustment of the rate matrix without retraining, addressing limitations of continuous-time diffusion graph models (Siraudin et al., 2024; Xu et al., 2024). This decoupling supports extensive, training-free performance optimization during the sampling stage, which is crucial for improving performance and reducing the number of sampling steps. Below, we propose the key components of DeFoG that are enabled by this disentanglement.
Sample Distortion
In DFM’s sampling process, the discretization is performed using equally sized time steps (Alg. 2, line 5). However, this uniformity may fail to preserve key properties during critical intervals where finer control is needed. For instance, smaller steps are essential near the final stages of sampling to prevent sudden edge alterations that could compromise global properties such as planarity. To overcome this limitation, we propose using variable step sizes, allocating smaller, more frequent steps during these critical intervals to better capture essential graph characteristics. Specifically, we apply to each timestep $t$ a bijective, increasing distortion function $f$ defined for $t\in[0,1]$, yielding $t^{\prime}=f(t)$. For example, the choice $f(t)=2t-t^{2}$ (referred to as polydec) creates monotonically decreasing step sizes, emphasizing the final stages of sampling, where error correction can be most critical. The specific distortion functions employed are described in Sec. B.1. Importantly, we can efficiently (i.e., without re-training) adjust the sample distortion adopted for each dataset to better accommodate its graphs characteristics, leading to significant performance improvements.
Train Distortion
Once the optimal sampling distortion is identified, it can guide training by highlighting the critical time ranges for graph generation in a specific dataset. This enables adjustments to the training distribution $\mathcal{T}$, skewing it toward these ranges to focus the model’s capacity on the most relevant regions. The skewed distributions are obtained by passing uniformly sampled times $t$ through the same distortion functions. While similar strategies in other modalities, such as image generation (Esser et al., 2024), often emphasize intermediate time ranges, we find that optimal time ranges in graph generation vary across datasets. Aligning the distortion function in training with sampling typically enhances the algorithm by focusing on critical time ranges. For instance, for larger datasets, such as drug-sized molecular graphs, the polydec distortion function accelerates convergence significantly and provides noticeable performance improvements.
Target Guidance
The application of time distortions is not the sole avenue for optimizing the sampling process; the design of the conditional rate matrices also offers significant potential for improvement. One promising direction arises from the goal of better guiding the generation process toward the clean data distribution (Song et al., 2020a). This also aligns with the fundamental design of diffusion and flow matching models, which are structured to predict clean data directly and subsequently use that prediction to generate the denoising trajectory (Ho et al., 2020; Lipman et al., 2023; Vignac et al., 2022). Inspired by these principles, we propose an alternative sampling mechanism that seeks to further amplify the influence of the denoising neural network’s predictions in the designed rate matrices, by setting $R_{t}{(z_{t},z_{t+\mathrm{d}t}|z_{1})}=R^{*}_{t}{(z_{t},z_{t+\mathrm{d}t}|z_{1})}+R^{\omega}_{t}(z_{t},z_{t+\text{d}t}|z_{1})$ for $z_{t}\neq z_{t+\mathrm{d}t}$, such that:
$R^{\omega}_{t}(z_{t},z_{t+\text{d}t}|z_{1})=\omega\frac{\delta(z_{t+\mathrm{d}t},z_{1})}{\mathcal{Z}^{>0}_{t}\;p_{t|1}(z_{t}|z_{1})}.\vspace{-3pt}$ (6)
This adjustment increases the weight of transitions in the rate matrix when $z_{t+\text{d}t}=z_{1}$, where $z_{1}$ is the predicted clean data. While moderate increases in $\omega$ significantly enhances performance by steering the generation toward high confidence domains, excessively high $\omega$ leads to performance drop. This behavior is explained in Lemma 10, in Sec. B.2, where we show that target guidance introduces an $O(\omega)$ violation of the Kolmogorov equation. Consequently, finding an optimal value for $\omega$ is essential.
Stochasticity
Orthogonal to target guidance, there also exists unexplored potential in the design space of conditional rate matrices that preserve the Kolmogorov equation as the standard formulation of $R^{*}_{t}{(z_{t},z_{t+\mathrm{d}t}|z_{1})}$ does not fully capture this space. As demonstrated by Campbell et al. (2024), for any rate matrix $R^{\text{DB}}_{t}$ satisfying the detailed balance condition $p_{t|1}(z_{t}|z_{1})R^{\text{DB}}_{t}(z_{t},z_{t+\text{d}t}|z_{1})=p_{t|1}(z_{t+\text{d}t}|z_{1})R^{\text{DB}}_{t}(z_{t+\text{d}t},z_{t}|z_{1})$, the modified rate matrix $R^{\eta}_{t}=R^{*}_{t}+\eta R^{\text{DB}}_{t}$, with $\eta\in\mathbb{R}^{+}$, is also valid. Intuitively, increasing $\eta$ facilitates more transitions to other states while reducing the likelihood of remaining in the same state, thereby increasing stochasticity in the trajectory of the denoising process. This approach can be interpreted as a correction mechanism, as it draws transitions back to states that would otherwise be forbidden according to the rate matrix formulation, as described in Sec. B.5. Additionally, different designs of $R^{\text{DB}}_{t}$ encode different priors for preferred transitions between states, which we investigate in detail in Sec. B.3.
3.3 Permutation Invariance Guarantees
Graph generative models should respect the inherent permutation symmetries of graphs. Accordingly, DeFoG’s training and sampling algorithms should be independent of node ordering. This requires that both DeFoG’s loss function during training and its probability of generating a specific graph during sampling be permutation invariant. We formally demonstrate these results in Lemma 3, whose proof is in Sec. D.2.1.
For any permutation equivariant denoising neural network, the loss function of DeFoG is permutation invariant, and its sampling probability is permutation invariant.
We further describe the permutation equivariant denoising neural network of DeFoG in Sec. 5.